# Enhancement of axial anomaly effects in hot two-color QCD: FRG approach in the linear sigma model

### v1

@@ III.2 General structures
Since chiral symmetry is realized linearly in our framework, the effective action (and thus the effective potential) inherits the symmetry properties of the classical action. That is, $V_{k}$ can only depend on independent chiral invariants, $V_{k}=V_{k}(I_{1},I_{2},I_{A})$. Within the present LSM framework, the spontaneous breaking of chiral symmetry is triggered by the appearance of a nonzero ${\cal S}^{0}$ mean field, $\sigma_{0}\equiv\langle{\cal S}^{0}\rangle$. In this vacuum configuration the $I_{2}$ invariant vanishes: $\langle I_{2}\rangle=0$, and hence, even after taking into account quantum corrections to the effective potential, higher orders of $I_{2}$, i.e., $I_{2}^{2}$, $I_{2}^{3}$, $\cdots$, are expected to produce small contributions. For this reason we employ the following ansatz for the potential at the quantum level:
@@
where $m_{k}^{2}$, $\lambda_{1,k}$, $a_{k}$, $c_{1,k}$, $c_{2,k}$ and $\lambda_{2,k}$ are field independent flowing coupling constants. Note that the structure of Eq. (31) is unchanged from the classical one (22), meaning that only renormalizable operators are taken into account. Each coupling explicitly contains a subscript $k$ stressing that they are defined at scale $k$.
@@
We note that both $({\cal D}_{k}^{-1})^{AB}$ and $(V_{k}^{\prime\prime})^{AB}$ are $12\times 12$ matrices since in total $12$ hadrons are considered. Using the above notation, from Eq. (III.1) we have
@@ III.3 Flowing couplings
The first term of Eq. (LABEL:OmegaExpanded) corresponds to the vacuum energy, which has nothing to do with the flow equations of the couplings. It is the second and third terms, which determine the latter in the end.
@@
which leads to flow equations for $m_{k}$ and $a_{k}$. The trace of Eq. (39) is readily performed since only one ${\cal D}_{0,k}$ is included. With the help of the Matsubara-summation formula
@@
We note that $E_{k}^{+}$ corresponds to $\sigma$, $\pi$, and $B$ ($\bar{B}$), while $E_{k}^{-}$ for $\eta$, $a_{0}$, $B^{\prime}$ ($\bar{B}^{\prime}$).
Since $\partial_{k}V_{M,k}^{(2)}$ and $\partial_{k}V_{B,k}^{(2)}$ are proportional to quadratic forms in terms of ${\cal S}^{a}$ and ${\cal P}^{a}$, due to symmetry they must combine into contributions proportional to $I_{1}$ and $I_{A}$ appropriately, i.e.,
$\partial_{k}V_{M,k}^{(2)}$ $=$ $\partial_{k}\hat{V}^{I_{1}}_{M,k}I_{1}+\partial_{k}\hat{V}^{I_{A}}_{M,k}I_{A}\ ,$
$\partial_{k}V_{B,k}^{(2)}$ $=$ $\partial_{k}\hat{V}^{I_{1}}_{B,k}I_{1}+\partial_{k}\hat{V}^{I_{A}}_{B,k}I_{A}\ .$ (45)
Inserting (III.3) into (41) yields
@@
Therefore, the flow equations for $m_{k}$ and $a_{k}$ can be easily obtained as
$\partial_{k}m_{k}$ $=$ $\partial_{k}\hat{V}^{I_{1}}_{M,k}+\partial_{k}\hat{V}^{I_{1}}_{B,k}\ ,$
@@
and we note once again that, the ~$\tilde{\partial}_{k}$ operator acts only on $R_{k}$ in Eqs. (48) - (50). We stress that in Eq. (51), mesonic and baryonic contributions are intentionally separated, as it turns out to be useful for later analysis. The ~$\tilde{\partial}_{k}$ operation in the RHSs is easily handled by means of Eq. (A). Note that, since $\partial_{k}R_{k}({\bm{p}})\sim\Theta(k^{2}-{\bm{p}}^{2})$, all momentum integrals are restricted to $|{\bm{p}}|\in[0,k]$, leading to ${\bm{p}}^{2}+R_{k}({\bm{p}})\rightarrow k^{2}$ in the integrands. As a result, all integrals become trivial, and one formally substitutes $\int d^{3}p/(2\pi)^{3}\rightarrow k^{3}/6\pi^{2}$.
Unlike demonstrated above, the third term in the RHS of Eq. (39) contains three types of contributions, i.e., purely mesonic ones, purely baryonic ones, and those that mix the two. In accordance, in the forthcoming analysis, we define the first one as the purely mesonic contribution, while the second and third ones will belong to the contributions classified as baryonic. This separation enables us to obtain flow equations for $\lambda_{1,k}$, $\lambda_{2,k}$, $c_{1,k}$ and $c_{2,k}$ as a sum of mesonic and baryonic fluctuations. We do not go through the explicit calculations for the aforementioned flows, they are summarized in Appendix. A.
Before closing this subsection, we comment on the anomalous structure of the effective potential. When the $U(1)_{A}$ anomaly is switched off, $a_{k}=0=c_{1,k}=c_{2,k}$ hold. In this case the $U(1)$ axial charges are conserved, or intuitively speaking, chirality-flipping processes of a single quark do not occur. As a result, all fluctuations universally obey the dispersion relation of $E_{k}^{+}=E_{k}^{-}=\sqrt{{\bm{p}}^{2}+R_{k}({\bm{p}})+m_{k}^{2}}$. Consequently, all the anomalous components of the potential vanish: $\hat{V}^{I_{A}}_{M,k}=\hat{V}^{I_{A}}_{B,k}=\hat{V}_{M,k}^{I_{A}^{2}}=\hat{V}_{B,k}^{I_{A}^{2}}=\hat{V}_{M,k}^{I_{1}I_{A}}=\hat{V}_{M,k}^{I_{1}I_{A}}=0$.
@@
$0.3$ $-0.0426$ $-0.032$ $1.25$ $5$ 00 00
$0.5$ $0.0465$ $-0.108$ $1.25$ $5$ 00 00
$0.95$ $0.385$ $-0.428$ $1.25$ $5$ 00 00
Table 1: The parameters at $k=\Lambda_{\rm UV}$ for $\mathring{M}_{\eta}=0.3$, $0.5$, $0.95$ GeV with $\Lambda_{\rm UV}=1$ GeV.
@@ IV Numerical results
In this section we proceed with numerical analyses by solving the flow equations derived in Sec. III.3.
First, in Sec. IV.1 our criterion to fix initial conditions for the flow equations is explained. Then, in Sec. IV.2 and Sec. IV.3 we present our main results on the hadron masses and modifications of the anomaly coefficients at finite temperature. Finally, in Sec. IV.4, numerical results on the topological susceptibility are shown.
@@ IV.1 Inputs
We have six parameters in the potential (22), which also correspond to initial conditions of the flow equations. In the present analysis, we assume the large $N_{c}$ ansatz for the potential for the UV potential, where the double-trace terms and the anomalous four-point interactions are suppressed Witten (1979); Di Vecchia and Veneziano (1980), leading to $\lambda_{2,\Lambda}=4\lambda_{1,\Lambda}$ [see Eq. (1)] and $c_{1,\Lambda}=c_{2,\Lambda}=0$. In addition, it is found in the numerics that $\lambda_{2,\Lambda}$ must satisfy $\lambda_{2,\Lambda}\lesssim 5$ to produce a reasonable reduction of the chiral condensate at finite temperature. Hence, we adopt $\lambda_{2,\Lambda}=5$ and accordingly $\lambda_{1,\Lambda}=1.25$ as an adequate demonstration. The quadratic couplings at at the UV scale, $m_{\Lambda}^{2}$ and $a_{\Lambda}$ are then fixed by fitting the pion and the $\eta$ meson masses in the vacuum. The former will be fixed to $\mathring{M}_{\pi}=0.14$ GeV, while for the latter we investigate three different scenarios, $\mathring{M}_{\eta}=0.3,0.5,0.95$ GeV [^3] The symbol $\mathring{M}$ stands for the corresponding vacuum value. Also, in what follows, we will omit the subscript 00 representing $k=0$ in the masses and parameters, as a shorthand notation.. The reason is that our main motivation in this study is to explore the fate of $U(1)_{A}$ anomaly at finite temperature, and one might be curious about such dependencies as a function of the anomaly strength itself. All the masses are evaluated at $T=0$, after integrating down the flow equations to $k=0$, see their respective expressions, (75). Finally, the explicit breaking parameter $\bar{c}m_{q}$ in Eq. (20) is determined by the decay constant, $f_{\pi}=0.093$ GeV, which equals to the vacuum expectation value $\mathring{\sigma}_{0}$ due to the corresponding Ward identity. In our study, the initial UV scale is set to be $\Lambda_{\rm UV}=1$ GeV, and all the numerical values of the determined parameters are summarized in Table 1.
Figure 1: $T$ dependencies of the masses of $\eta$, $\pi$, $\sigma$ and $a_{0}$ mesons with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV.
Figure 2: $T$ dependencies of the mass differences $\Delta M_{(\sigma.\pi)}$ and $\Delta M_{(a_{0}.\eta)}$ with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV.
@@ IV.2 Hadron masses at finite $T$
In this subsection, we examine $T$ dependencies of the hadron masses by solving the flow equations. Their analytic mass formulas are summarized in Appendix B, which indicate that $M_{\pi}=M_{B(\bar{B})}$ while $M_{a_{0}}=M_{B^{\prime}(\bar{B}^{\prime})}$ at any $T$, stemming from the pseudoreality ${\bm{2}}\simeq\bar{\bm{2}}$ of QC2Suenaga et al. (2023); Suenaga (2025).
Figure 1 shows the resultant hadron masses for $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV at finite $T$. All the masses are mostly dominated by the quadratic terms $m^{2}$ and $a$, since the quartic couplings are suppressed, particularly $\lambda_{2}$ (and accordingly $\lambda_{1}=\lambda_{2}/4$) are found to remain small throughout the flow. Hence, from the mass formulas of the hadrons provided in Eqs. (75) and (LABEL:SigmaMass), mass differences between $\sigma$ and $\pi$ or between $a_{0}$ and $\eta$ are always small as $M^{2}_{\sigma},M^{2}_{\pi}\sim m^{2}+a$ while $M^{2}_{a_{0}},M^{2}_{\eta}\sim m^{2}-a$. We note that this approximation applies to any $k$ throughout the flow.
Within the present approximation, the strength of thermal excitations is determined by the size of the effective masses (44); smaller values of $m_{{\rm eff},k}^{\pm}$ throughout the flows yield more thermal effects in the end. Thus, from the above argument one can understand that a smaller value of $\mathring{M}_{\eta}$ results in more significant thermal contributions to the potential. This property is also reflected by the reduction rate of the chiral condensate, as shown in Fig. 4. In fact, this figure indicates that the smaller $\mathring{M}_{\eta}$ we take, the more rapidly the chiral condensate drops.
Figure 3: $T$ dependencies of the normalized anomaly coefficients $\tilde{a}$, $\tilde{c}_{1}$ and $\tilde{c}_{2}$ defined in Eq. (53), with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV.
@@
in Fig. 2. At finite $T$, the mass differences reduce because of chiral symmetry restoration, for which the chiral partner structures are also clearly realized. Similar mass degeneracies were also found at the high density limit with vanishing temperature Suenaga et al. (2023, 2024). We note that $\Delta M_{(\sigma,\pi)}$ and $\Delta M_{(a_{0},\eta)}$ are at most of ${\cal O}$(10 MeV), regardless of the value of $\mathring{M}_{\eta}$, since the $\eta$ meson mass is predominantly generated by the axial anomaly, as shown in Eq. (80), and the anomaly does not play a significant role in the mass degeneracies of chiral partners.

### v2

@@ III.2 General structures
Since chiral symmetry is realized linearly in our framework, the effective action (and, thus, the effective potential) inherits the symmetry properties of the classical action. That is, $V_{k}$ can depend on only independent chiral invariants, $V_{k}=V_{k}(I_{1},I_{2},I_{A})$. Within the present LSM framework, the spontaneous breaking of chiral symmetry is triggered by the appearance of a nonzero ${\cal S}^{0}$ mean field, $\sigma_{0}\equiv\langle{\cal S}^{0}\rangle$. In this vacuum configuration the $I_{2}$ invariant vanishes: $\langle I_{2}\rangle=0$. After taking into account quantum corrections to the effective potential, higher orders of $I_{2}$, i.e., $I_{2}^{2}$, $I_{2}^{3}$,…, are expected to be generated. These are, however, nonrenormalizable interactions, which are expected to produce only small contributions during the flows. For this reason, we employ the following ansatz for the potential at the quantum level:
@@
where $m_{k}^{2}$, $\lambda_{1,k}$, $a_{k}$, $c_{1,k}$, $c_{2,k}$ and $\lambda_{2,k}$ are field-independent flowing coupling constants. Note that the structure of Eq. (32) is unchanged from the classical one (22), meaning that only renormalizable operators are taken into account. Each coupling explicitly contains a subscript $k$ stressing that they are defined at scale $k$.
@@
We note that both $({\cal D}_{k}^{-1})^{AB}$ and $(V_{k}^{\prime\prime})^{AB}$ are $12\times 12$ matrices. since in total $12$ hadrons are considered. Using the above notation, from Eq. (III.1) we have
@@ III.3 Flowing couplings
The first term in Eq. (LABEL:OmegaExpanded) corresponds to the vacuum energy, which has nothing to do with the flow equations of the couplings. It is the second and third terms which determine the latter in the end.
@@
which leads to flow equations for $m_{k}$ and $a_{k}$. The trace of Eq. (40) is readily performed, since only one ${\cal D}_{0,k}$ is included. With the help of the Matsubara-summation formula
@@
We note that $E_{k}^{+}$ corresponds to $\sigma$, $\pi$, and $B$ ($\bar{B}$), while $E_{k}^{-}$ for $\eta$, $a_{0}$, and $B^{\prime}$ ($\bar{B}^{\prime}$).
Since $\partial_{k}V_{M,k}^{(2)}$ and $\partial_{k}V_{B,k}^{(2)}$ are proportional to quadratic forms in terms of ${\cal S}^{a}$ and ${\cal P}^{a}$, due to symmetry their sum must combine into contributions proportional to $I_{1}$ and $I_{A}$ appropriately, i.e.,
$\partial_{k}V_{M,k}^{(2)}$ $=$ $\partial_{k}\hat{V}^{I_{1}}_{M,k}I_{1}+\partial_{k}\hat{V}^{I_{A}}_{M,k}I_{A}+\partial_{k}\delta V_{45,k}\ ,$
$\partial_{k}V_{B,k}^{(2)}$ $=$ $\partial_{k}\hat{V}^{I_{1}}_{B,k}I_{1}+\partial_{k}\hat{V}^{I_{A}}_{B,k}I_{A}-\partial_{k}\delta V_{45,k}\ ,$ (46)
where $\delta V_{45,k}$ is a function of $P^{4}$, $P^{5}$, $S^{4}$ and $S^{5}$. This $\delta V_{45,k}$ appears to ensure the $SU(4)$-symmetry property, which relates mesons and baryons, and it drops out from $\partial_{k}V_{k}^{(2)}$ in Eq. (42). Here, we only stress that when considering mesonic and baryonic contributions separately in Sec. V, we will omit $\delta V_{45,k}$. Inserting (III.3) into (42) yields
@@
Therefore, the flow equations for $m^{2}_{k}$ and $a_{k}$ can be easily obtained as
$\partial_{k}m^{2}_{k}$ $=$ $\partial_{k}\hat{V}^{I_{1}}_{M,k}+\partial_{k}\hat{V}^{I_{1}}_{B,k}\ ,$
@@
and we note once again that the $\tilde{\partial}_{k}$ operator acts only on $R_{k}$ in Eqs. (48)–(51). We stress that, in Eq. (52), mesonic and baryonic contributions are intentionally separated, as it turns out to be useful for later analysis. The $\tilde{\partial}_{k}$ operation in the rhs is easily handled by means of Eq. (A). Note that, since $\partial_{k}R_{k}({\bm{p}})\sim\Theta(k^{2}-{\bm{p}}^{2})$, all momentum integrals are restricted to $|{\bm{p}}|\in[0,k]$, leading to ${\bm{p}}^{2}+R_{k}({\bm{p}})\rightarrow k^{2}$ in the integrands. As a result, all integrals become trivial, and one formally substitutes $\int d^{3}p/(2\pi)^{3}\rightarrow k^{3}/6\pi^{2}$.
Unlike demonstrated above, the third term in the rhs of Eq. (LABEL:OmegaExpanded) contains three types of contributions, i.e., purely mesonic ones, purely baryonic ones, and those that mix the two. In accordance, in the forthcoming analysis, we define the first one as the purely mesonic contribution, while the second and third ones will belong to the contributions classified as baryonic. This separation enables us to obtain flow equations for $\lambda_{1,k}$, $\lambda_{2,k}$, $c_{1,k}$, and $c_{2,k}$ as a sum of mesonic and baryonic fluctuations. We do not go through the explicit calculations for the aforementioned flows; they are summarized in Appendix A.
Before closing this subsection, we comment on the anomalous structure of the effective potential. When the $U(1)_{A}$ anomaly is switched off, $a_{k}=0=c_{1,k}=c_{2,k}$ holds. In this case, the $U(1)$ axial charges are conserved, or, intuitively speaking, chirality-flipping processes of a single quark do not occur. As a result, all fluctuations universally obey the dispersion relation of $E_{k}^{+}=E_{k}^{-}=\sqrt{{\bm{p}}^{2}+R_{k}({\bm{p}})+m_{k}^{2}}$. Consequently, all the anomalous components of the potential vanish: $\hat{V}^{I_{A}}_{M,k}=\hat{V}^{I_{A}}_{B,k}=\hat{V}_{M,k}^{I_{A}^{2}}=\hat{V}_{B,k}^{I_{A}^{2}}=\hat{V}_{M,k}^{I_{1}I_{A}}=\hat{V}_{M,k}^{I_{1}I_{A}}=0$.
@@
$0.3$ 0.0426-0.0426 0.032-0.032 $1.25$ 55 0 0
$0.5$ $0.0465$ 0.108-0.108 $1.25$ 55 0 0
$0.95$ $0.385$ 0.428-0.428 $1.25$ 55 0 0
Table 1: The parameters at $k=\Lambda_{\rm UV}$ for $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV with $\Lambda_{\rm UV}=1$ GeV.
@@ IV Numerical results
In this section, we proceed with numerical analyses by solving the flow equations derived in Sec. III.3.
First, in Sec. IV.1, our criterion to fix initial conditions for the flow equations is explained. Then, in Secs. IV.2 and Sec. IV.3, we present our main results on the hadron masses and modifications of the anomaly coefficients at finite temperature. Finally, in Sec. IV.4, numerical results on the topological susceptibility are shown.
Figure 1: $T$ dependencies of the masses of $\eta$, $\pi$, $\sigma$, and $a_{0}$ mesons with $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV.
Figure 2: $T$ dependencies of the mass differences $\Delta M_{(\sigma.\pi)}$ and $\Delta M_{(a_{0}.\eta)}$ with $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV.
Figure 3: $T$ dependencies of the normalized anomaly coefficients $\tilde{a}$, $\tilde{c}_{1}$, and $\tilde{c}_{2}$ defined in Eq. (54), with $\mathring{M}_{\eta}=0.3$, $0.5$ and $0.95$ GeV.
@@ IV.1 Inputs
We have six parameters in the potential (22), which also correspond to initial conditions of the flow equations. In the present analysis, we assume the large $N_{c}$ ansatz for the potential for the UV potential, where the double-trace terms and the anomalous four-point interactions are suppressed Witten:1979kh; DiVecchia:1980yfw, leading to $\lambda_{2,\Lambda}=4\lambda_{1,\Lambda}$ [see Eq. (23)] and $c_{1,\Lambda}=c_{2,\Lambda}=0$. In addition, it is found in the numerics that $\lambda_{2,\Lambda}$ must satisfy $\lambda_{2,\Lambda}\lesssim 5$ to produce a reasonable reduction of the chiral condensate at finite temperature. [^2] When choosing $\lambda_{2}\gtrsim 5$, the chiral condensate can increase at finite $T$, which is obviously unphysical. Hence, we adopt $\lambda_{2,\Lambda}=5$ and accordingly $\lambda_{1,\Lambda}=1.25$ as an adequate demonstration. The quadratic couplings at at the UV scale, $m_{\Lambda}^{2}$ and $a_{\Lambda}$, are then fixed by fitting the pion and the $\eta$ meson masses in the vacuum. The former will be fixed to $\mathring{M}_{\pi}=0.14$ GeV, while for the latter we investigate three different scenarios, $\mathring{M}_{\eta}=0.3,0.5,0.95$ GeV. [^3] The symbol $\mathring{M}$ stands for the corresponding vacuum value. Also, in what follows, we will omit the subscript 0 representing $k=0$ in the masses and parameters, as a shorthand notation. The reason is that our main motivation in this study is to explore the fate of $U(1)_{A}$ anomaly at finite temperature, and one might be curious about such dependencies as a function of the anomaly strength itself. All the masses are evaluated at $T=0$, after integrating down the flow equations to $k=0$: see their respective expressions, (79). Finally, the explicit breaking parameter $\bar{c}m_{q}$ in Eq. (20) is determined by the decay constant $f_{\pi}=0.093$ GeV, which equals to the vacuum expectation value $\mathring{\sigma}_{0}$ due to the corresponding Ward identity. In our study, the initial UV scale is set to be $\Lambda_{\rm UV}=1$ GeV, and all the numerical values of the determined parameters are summarized in Table 1. We also note that current lattice simulations in QC2D are being carried out with a heavy pion mass, e.g., $\mathring{M}_{\pi}^{\rm lattice}\sim 0.7$ GeV Boz:2019enj; Astrakhantsev:2020tdl; Iida:2024irv. Hence, predictions of our present work may also be considered as an important benchmark for future lattice computations with lighter pion masses.
@@ IV.2 Hadron masses at finite $T$
In this subsection, we examine $T$ dependencies of the hadron masses by solving the flow equations. Their analytic mass formulas are summarized in Appendix B, which indicate that $M_{\pi}=M_{B(\bar{B})}$ while $M_{a_{0}}=M_{B^{\prime}(\bar{B}^{\prime})}$ at any $T$, stemming from the pseudoreality ${\bm{2}}\simeq\bar{\bm{2}}$ of QC2Suenaga:2022uqn; Suenaga:2025sln.
Figure  1 shows the resultant hadron masses for $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV at finite $T$. All the masses are mostly dominated by the quadratic terms $m^{2}$ and $a$, since the quartic couplings are suppressed; particularly, $\lambda_{2}$ (and, accordingly, $\lambda_{1}=\lambda_{2}/4$) are found to remain small throughout the flow. Hence, from the mass formulas of the hadrons provided in Eqs. (79) and (LABEL:SigmaMass), mass differences between $\sigma$ and $\pi$ or between $a_{0}$ and $\eta$ are always small as $M^{2}_{\sigma},M^{2}_{\pi}\sim m^{2}+a$ while $M^{2}_{a_{0}},M^{2}_{\eta}\sim m^{2}-a$. We note that this approximation applies to any $k$ throughout the flow.
Within the present approximation, the strength of thermal excitations is determined by the size of the effective masses (45); smaller values of $m_{{\rm eff},k}^{\pm}$ throughout the flows yield more thermal effects in the end. Thus, from the above argument one can understand that a smaller value of $\mathring{M}_{\eta}$ results in more significant thermal contributions to the potential. This property is also reflected by the reduction rate of the chiral condensate, as shown in Fig. 4. In fact, this figure indicates that the smaller $\mathring{M}_{\eta}$ we take, the more rapidly the chiral condensate drops. The pseudocritical temperature read off from Fig. 4 is $T_{\rm pc}\gtrsim 0.25$ GeV, which is higher than the one for $N_{c}=3$, where $T_{\rm pc}$ is estimated to be $T_{\rm pc}\sim 0.15$ GeV. The slightly higher $T_{\rm pc}$ is also our prediction for QC2D with a light pion mass.
Figure 1 indicates that the $\sigma$ meson mass is always $\mathring{M}_{\sigma}\sim 0.18$ GeV at vanishing temperature, which is smaller than the two-pion threshold $2\mathring{M}_{\pi}=0.28$ GeV. That is, our current analysis predicts a stable $\sigma$ meson in QC2D with the physical pion mass, unlike the $N_{c}=3$ world. It would be interesting to see whether future lattice simulations can confirm such a light $\sigma$ meson.
@@
in Fig. 2. At finite $T$, the mass differences reduce because of chiral symmetry restoration, for which the chiral partner structures are also clearly realized. Similar mass degeneracies were also found at the high-density limit with vanishing temperature Suenaga:2022uqn; Suenaga:2023xwa. We note that $\Delta M_{(\sigma,\pi)}$ and $\Delta M_{(a_{0},\eta)}$ are at most of ${\cal O}(10)$ MeV, regardless of the value of $\mathring{M}_{\eta}$, since the $\eta$ meson mass is predominantly generated by the axial anomaly, as shown in Eq. (84), and the anomaly does not play a significant role in the mass degeneracies of chiral partners. The mass difference $\Delta M_{(\sigma,\pi)}$ gets enhanced for a heavier $\eta$ meson, which mainly results from the increase of $c_{2}$. We note that, when the anomaly is turned off, the tendency is opposite; the mass difference $\Delta M_{(\sigma,\pi)}$ is reduced for heavier $\mathring{M}_{\eta}$ when $a=c_{1}=c_{2}=0$.
Figure 4: $T$ dependencies of the mean field $\sigma_{0}$ with $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV.
Figure 5: $T$ dependencies of the normalized topological susceptibilities $\chi_{\rm top}$ (solid lines) and $\tilde{\chi}_{\rm top}$ (dashed lines) defined in Eqs. (58) and (59), respectively, with $\mathring{M}_{\eta}=0.3$, $0.5$, and $0.95$ GeV.